OpenStudy (anonymous):

use the function to answer the question: x+14 if x<-7 f(x) 21 if -7 ≤x<-6 x^2 -1 if x ≥ -6 what is the value of f(-6)?

OpenStudy (anonymous):

Since u want to find F(x) at x=-6 , U will use the Function F(x)= x^2-1 because x is bigger or **Equal** -6 Plug -6 in the equation so it will be F(-6)=(-6)^2-1=.....

OpenStudy (anonymous):

35?

hartnn (hartnn):

yup,35

OpenStudy (anonymous):

Bingo!

OpenStudy (anonymous):

what about f(x)=x^2-4; g(x)=x+2 which is equal to (f+g)(x)

OpenStudy (anonymous):

U mean Find F(F(x)+g(x)) ?

OpenStudy (anonymous):

no its f(x)+g(x) im not to sure but i think that might be it

hartnn (hartnn):

then just add them x^2-4+x+2 = ?

OpenStudy (anonymous):

x^2+x-2?

hartnn (hartnn):

yup

OpenStudy (anonymous):

okay so what is f(x)=4x-1 g(x)=5x^2 which is equal to (fºg)(x)?

hartnn (hartnn):

for (fog)(x) , replace 'x' in f(x) by g(x) = 5x^2

OpenStudy (anonymous):

so then : 4(5x^2)-1

hartnn (hartnn):

yup, that will be 20x^2-1

OpenStudy (anonymous):

and then what is f(x)=4x-7 g(x)=x+3 what is the function of (gºf)(4)

OpenStudy (anonymous):

is it 12?

hartnn (hartnn):

to get (gof)(x) replace 'x' in g(x) by f(x) and then put x=4

hartnn (hartnn):

yup, its 12

OpenStudy (anonymous):

what is f(x)=5x-10 what is the value of f^-1 (-3)

hartnn (hartnn):

so got that, its 1.4

OpenStudy (anonymous):

yea

hartnn (hartnn):

any more questions to come ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what is the vertex of the graph of f(x)=-2|x-5|+3

OpenStudy (anonymous):

i think its -5, 3

OpenStudy (anonymous):

+5,3

hartnn (hartnn):

nopes, (5,3) vertex for f(x) = p|x+a|+b is -a,b

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

the graph of g(x) results when the graph of f(x)=-|x| is shifted 3 units to the left. i think it's g(x)=-|x+3|

OpenStudy (anonymous):

what is the value of f(-4) when f(x)= -10 if x< -5 x^3 if -5≤x≤2 2x+4 if x>2

hartnn (hartnn):

in which range does -4 lie ?

OpenStudy (anonymous):

the middle one and the last one

hartnn (hartnn):

not the last one, x>2 means x cannot be -4 so only middle one. put x= -4 in middle function, what do u get ?

OpenStudy (anonymous):

-64

hartnn (hartnn):

thats your answer :)

OpenStudy (anonymous):

okay now this one is kinda confusing to me but i think i know the answer....

OpenStudy (anonymous):

i think it is the third one, but im not too sure....

hartnn (hartnn):

nopes, its 4th. put these values, x=0,0.5,1,1.5,..., find g(x) u will get it

OpenStudy (anonymous):

ohhhh i get it!!! its soo confusing sometimes!

OpenStudy (anonymous):

f(x)=2x+5 g(x)=3x^2 which is equal to (fºg)(x)? i tried it and i got 3x^2+2x+5 but i think it's wrong....

hartnn (hartnn):

put x= 3x^2 in f(x) 2(3x^2)+5 = ?

OpenStudy (anonymous):

6x^2+5?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

omg you make it easier to understand hahaha

hartnn (hartnn):

more questions ? i need to go...

OpenStudy (anonymous):

yea one last one f(x)=2/x g(x)=|x|-1 what is the domain of (fºg)(x)

hartnn (hartnn):

fog (x) = ?

OpenStudy (anonymous):

x|x=real numbers x|x ≠0 x|x≠0, x≠1 x|x≠-1, x≠1

hartnn (hartnn):

f(g(x)) = 2/(|x|-1) ok? for which values is denominator = 0 ??

OpenStudy (anonymous):

no the ones i put up last were the options

hartnn (hartnn):

i know. denominator should not be 0 |x|-1 should not be 0 for x=1, |1|-1=1-1=0 for x=-1, |-1|-1=1-1=0 so x|x≠-1, x≠1 got it ?

OpenStudy (anonymous):

ohh okay yea! thank you! you make it easier to understand ty very much

hartnn (hartnn):

welcome :)